# Quantifying the Dunkelflaute: An analysis of variable renewable energy droughts in Europe

### v1

@@ 3 Discussion
Variable renewable energy droughts manifest in a wide range of patterns, from brief and isolated to very long-lasting, contiguous events of varying severity. Our multi-threshold analysis reveals that drought frequency, return periods and duration are highly sensitive to the chosen threshold. We further find that technology-specific characteristics are relevant for drought analyses, such as diurnal and seasonal patterns of solar PV in Europe. We thus argue that a multi-threshold approach is required to adequately characterize the frequency, duration, and severity of heterogeneous and sequential renewable energy drought patterns. Single-threshold analyses, which are prevalent in the literature, are not capable of detecting such patterns and lead to an incomplete characterization of extreme droughts.
@@
Based on a large data set and using a wide range of thresholds, we demonstrate that the complementarity of wind and solar power in Europe effectively reduces both the frequency and maximum duration of VRE droughts. Within isolated countries, combining solar and wind power leverages a portfolio effect, decreasing the maximum drought duration compared to single renewable technologies by on average 64% (PV), 52% (onshore wind), or 47% (offshore wind). This effect is particularly significant for PV, as wind power compensates for diurnal and seasonal shortages of solar energy. In the copperplate scenario, the portfolio effect for PV is even more substantial because the complementarity of wind power is more pronounced when considering a renewable portfolio covering a larger geographical area. We further show that an unconstrained geographical balancing between European countries gives rise to a substantial balancing effect, which shortens the longest PV, onshore wind, offshore wind, or technology portfolio droughts by 1%, 46%, or 34%, or 65% on average. This effect is driven by an imperfect spatio-temporal correlation of extreme drought periods, especially concerning wind droughts.
@@
To overcome the challenges of single-threshold analyses, we introduce a novel drought mass indicator that enables an integrated assessment of the duration and severity of events. We use it to identify extreme drought events of varying severity that may emerge sequentially within contiguous low-availability periods lasting several weeks or months. Such events may affect multiple European countries simultaneously, yet with varying intensity. In most countries and in a perfectly interconnected Europe, the most extreme drought events occur in winter. In wind-dominated systems, however, the most extreme events may occur during summer. We show that compound winter events, characterized by extreme technology portfolio droughts in combination with peak electricity demand, determine the major discharging period of long-duration electricity storage in a fully renewable European energy system.
@@
The severity of such events varies substantially across years and countries, enabling geographical balancing to mitigate the need for system flexibility. Using the drought mass metric, we find that the most severe renewable technology portfolio drought observable in the data occurred in the winter of 1996/97, assuming a perfectly interconnected, fully renewable European energy system. This European super drought event lasted 55 days. The maximum droughts in individual, isolated countries are much longer, e.g., 109 days in Germany. This shows that even during the most pronounced European drought event, geographical balancing can be leveraged.
@@
Our multi-threshold analysis facilitates a comprehensive characterization of renewable drought events in Europe, solely based on VRE availability time series. Our approach, and particularly the drought mass metric, is not prone to artifacts that may arise in drought analyses relying on energy system modeling such as the duration of the planning horizon or technology cost assumptions. These may substantially affect optimal long-duration storage investments [dowling_role_2020]. Yet, methodological challenges remain. Drought analyses purely based on availability time series cannot take into account real-world interconnection levels, which are between our extreme assumptions in the island and copperplate scenarios. Drought characteristics of scenarios with more policy-relevant interconnection levels are thus likely between these corner solutions. For instance, the duration of the European super drought is likely to last longer than the 55 days identified for the copperplate scenario, but shorter than identified here for isolated countries (Figure 7), depending on future interconnection capacities between countries.
@@
Due to computational limits, energy system models are often solved for a limited number of weather years, or even for only a single weather year. While substantial inter-annual variation in optimal energy model outcomes has been documented in the literature [pfenninger_long-term_2016, staffell_using_2016, pfenninger_dealing_2017, zeyringer_designing_2018, collins_impacts_2018, schlott_impact_2018, kaspar_climatological_2019, hill_effects_2021, grochowicz_intersecting_2023, richardson_climate_2023, gunn_spatial_2023, ruggles_planning_2024], our analysis shows that extreme portfolio droughts also vary significantly across different years and regions. This likely aggravates inter-annual variations of energy model outcomes, particularly in terms of long-duration system flexibility needs. Accordingly, it appears desirable to corroborate the findings and conclusions of previous model-based energy system studies that are based on a small number of weather years and do not include extreme VRE droughts. For example, the European Ten Year Network Development Plan 2022 draws on the single weather year 2009 [entso-e_tyndp_2022]. Long-term energy scenarios developed for the European Commission generally lack transparency about actual weather years used, but appear to be based on a very limited number of weather years [european_commission_directorate_general_for_energy_eu_2021]. The same holds true for influential climate neutrality scenarios for Germany [bdi_klimapfade_2021, fraunhofer_isi_langfristszenarien_2024, prognos_klimaneutrales_2021, luderer_deutschland_2021].
@@
We see several promising avenues for future research. We note that a wide range of different datasets of renewable availability are used for energy system modeling, with sometimes conflicting model outcomes and policy recommendations [kies_critical_2021]. Assessing the sensitivity of VRE drought characteristics to such datasets could provide valuable feedback to the meteorological research community that generates and continuously improves such datasets [craig_overcoming_2022]. Further, the temporal extent of most data sets is limited to a few decades. Statistically robust findings for low-probability VRE drought events with high impact are likely to require longer records of weather data, which also might involve a need for synthetic datasets, e.g., applied in Gangopadhyay et al. [gangopadhyay_role_2022]. Moreover, drought analyses for modeled future weather data that incorporate the effects of global climate change would be desirable. Finally, characterizing renewable energy droughts in other world regions and comparing these with our findings for Europe would be of interest. We argue that our multi-threshold framework is particularly useful for respective analyses in world regions where renewable availability is characterized by diurnal or seasonal variability, e.g., related to regional weather phenomena such as monsoon events.
@@ 4.1 Variable renewable energy drought definition and identification
Our VRE drought analysis is based on VRE availability factor time series. We use a new open-source method that applies a variable-duration mean below threshold (VMBT) algorithm for drought identification [kittel_measuring_2024]. This method searches for periods with a moving average below a given drought qualification threshold by iteratively decreasing the event duration. In each iteration, the algorithm sets the averaging interval to the respective event duration, starting with very long-lasting events and iteratively continuing to such that last only a few hours. A moving average below the drought threshold identifies a drought event. It is excluded from subsequent iterations, in which the averaging interval decreases further and additional (shorter) events are identified.
@@
This iterative procedure overcomes shortcomings of previous research [kittel_measuring_2024]. It pools adjacent periods that independently may not qualify as VRE drought, identifies unique events, avoids double counting as well as overlaps with adjacent events, and ensures that the full temporal extent of drought periods is captured [kittel_measuring_2024]. To cover the full spectrum of VRE droughts, we parameterize the VMBT algorithm such that it searches for periods ranging in duration from two years to one hour. The code and input data of the drought analysis tool are freely available on GitLab at https://gitlab.com/diw-evu/variable_renewable_energy_droughts_analyzer.
@@
To account for varying renewable generation potentials in the different settings, i.e., single technologies or different VRE portfolios across countries, we use a large number of drought thresholds and scale these relative to the multi-annual mean availability of each time series under investigation [kittel_measuring_2024]. Thresholds range from 10% to 100% of the mean availability factor across all investigated years, increasing in 5% increments. Very low thresholds around $\tau=0.1$ identify very severe droughts with near-zero VRE availability. In contrast, very high thresholds near $\tau=1.0$ are likely to identify below-average weather years. Figure SI.1 illustrates how these relative thresholds relate to absolute capacity factors.
@@
We use country-level VRE availability time series from the Pan-European Climate Database provided by the European Network of Transmission System Operators for Electricity [de_felice_entso-e_2022], comprising 38 weather years from 1982 to 2019. These data have been used for policy-relevant strategic reports, such as the Ten Year Network Development Plan (TYNDP) 2022 or the European Resource Adequacy Assessment 2021. The renewable capacity assumptions used to generate the capacity-weighted VRE portfolio time series stem from the TYNDP 2022 (scenario Distributed Energy). For Germany, we update the capacity targets according to the latest government targets [bundesamt_fur_justiz_gesetz_2024].
@@
4.2 Drought mass: a metric to identify storage-defining drought events
We introduce a drought mass metric to identify storage-defining drought events. For each investigated threshold $\tau\leq 0.75$, we use time series that include the hourly availability factors of identified events, assigning a value of 00 to hours that do not qualify as droughts according to a given threshold. We equally weigh across all thresholds by replacing the availability factors of identified events with the value $1$. For each hour, we then accumulate the resulting values. Within a given year-pair, we search for the highest cumulative multi-threshold score of contiguous events, using a cut-off threshold $\tau=0.75$. This approach defines the event duration according to the threshold $\tau=0.75$, while all thresholds $\tau\leq 0.75$ equally contribute to an event’s overall score. We tested a wide range of cut-offs and weighting schemes to identify the most effective drought mass design. The best alignment between events with the highest winter drought mass scores and major storage discharge events was achieved with a cut-off threshold of $\tau=0.75$ and equal weighting across all thresholds.
@@ 4.3 Power sector modeling and scenarios
We use the open-source power sector model Dispatch and Investment Evaluation Tool with Endogenous Renewables (DIETER) to analyze the interaction between VRE droughts and long-duration storage needs in a fully renewable European power sector. DIETER is a linear optimization model that determines least-cost capacity and dispatch decisions based on an hourly resolution [zerrahn2017, gaete-morales_dieterpy_2021]. Different versions of the model have been used to study various aspects of VRE integration and their interaction with other flexibility options or sector coupling technologies [zerrahn2018, schill_long-run_2018, schill_electricity_2020, schill_flexible_2020, kittel_renewable_2022, roth_geographical_2023, kirchem_power_2023, gong_bidirectional_2023, gaete-morales_power_2024]. Here, we use the most recent extension of the model that represents 33 European countries (EU27, the United Kingdom, Norway, Switzerland, and the Western Balkans). As very long-lasting VRE droughts may span across the turn of a calendar year, we extend DIETER’s planning horizon to two full years. For transparency and reproducibility, we provide the model code, the input data, and a manual in a public repository under permissive licenses available on GitLab at https://gitlab.com/diw-evu/projects/quantifying_the_dunkelflaute_energy_system_analysis.
@@ Acknowledgments
We thank the entire research group “Transformation of the Energy Economy” at the German Institute for Economic Research (DIW Berlin) for valuable inputs and discussions, as well as conference participants of the EGU General Assembly 2022, the International Energy Workshop 2022, the Conference on Climate, Weather and Carbon Risk in Energy and Finance 2022, the Next Generation Energy and Climate Workshop 2022, the International Conference Energy & Meteorology 2023, and the ENERDAY conference 2024 for valuable comments on earlier drafts. We acknowledge research grants by the Einstein Foundation (grant no. A-2020-612) and by the German Federal Ministry of Education and Research via the “Ariadne” projects (Fkz 03SFK5NO & 03SFK5NO-2).
@@
SI.2 Additional illustrations of frequency-duration distributions
@@
Figure SI.2 illustrates the cumulative frequency-duration distributions for droughts lasting up to a full year for Germany, Spain, and the European copperplate scenario. Only higher thresholds identify droughts lasting longer than a few weeks. For thresholds $\tau<1$, portfolio droughts are generally briefer than single-technology droughts. To raise complementary insights on seasonality, Figure SI.3 shows seasonally differentiated distributions for the copperplate scenario. In general, longer droughts are more frequent in winter than in summer. Brief PV droughts are more frequent in summer and are detected by higher thresholds.
@@
Figure SI.2: Exemplary cumulative frequency-duration distributions of drought events that may last up to one full year, sorting the frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific frequency.
@@
Figure SI.3: Exemplary cumulative frequency-duration distributions of drought events under the assumption of unconstrained geographical balancing, sorting the frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific frequency. For illustration, the distributions are truncated at 360 hours, i.e., they show events with a maximum duration of just above two weeks. Frequencies of events lasting at least two days, one week, and a fortnight are marked for a relative threshold $\tau=0.75$.
SI.3 Return periods: most extreme events occur rarely
@@
Return periods of VRE droughts are given by the reciprocal of yearly frequencies. The return period-duration distributions represent the right-hand side of the event distribution and indicate the longest event duration that can be expected to reoccur after a given number of years (Figure SI.4). Note that the return period monotonically increases and the maximum return period is limited by the 38 years of data that we investigate.
@@
For each threshold, the maximum event duration increases with higher return periods (compare increasing contour lines and marked data points in Figure SI.4). Additionally, the return period-duration distributions are highly sensitive to the underlying threshold. For any given return period, higher thresholds lead to events with longer maximum durations, plateauing at 365 days for very high thresholds. Note that such high durations indicate below-average wind and solar years and should not be mistaken for actual drought events. The return period-duration distributions vary significantly between technologies and countries. For example, the maximum offshore wind drought for a threshold of $\tau=0.75$ in Germany that returns every 20 years lasts 198 days and is much shorter than the corresponding PV (289 days) and onshore wind droughts (297 days). By comparison, in Spain, such 20-year return period droughts are shorter for PV (216 days) and onshore wind (245 days) but longer for offshore wind (211 days).
@@
When combined in a technology portfolio, the maximum duration decreases substantially for each return period in both Germany and Spain and also in the European copperplate, which is another instance of the portfolio effect described above. The balancing effect is also visible: with perfect interconnection, the maximum drought duration further decreases due to a limited temporal correlation of wind droughts across Europe. In the case of solar PV, higher solar availability in South Europe during winter balances lower availability in Northern Europe.
@@
Figure SI.4: Exemplary return period-duration distributions of rare drought events with an average annual frequency below 1. The contour lines along the return period axis represent the threshold-specific return period. The maximum duration of an event returning on average every 10, 20, and 30 years are marked for a threshold of $\tau=0.75$.
SI.4 Additional illustrations of maximum drought durations
@@
Figures SI.5 and SI.6 show the maximum duration of single drought events across all years for each country and threshold. Figure SI.7 illustrates the inter-annual variability of the maximum drought duration for each threshold and technology (portfolio).
@@
Figure SI.5: Most extreme duration of single drought event of all years for each country and threshold. The latter are sorted in descending order from left to right according to the longest duration for a threshold $\tau=0.75$.

### v2

@@ 3 Discussion
Variable renewable energy droughts manifest in a wide range of patterns, from brief and isolated to very long-lasting, contiguous events of varying severity. Our multi-threshold analysis reveals that drought frequency, return periods, and duration are highly sensitive to the chosen threshold. We thus argue that a multi-threshold approach is required to adequately characterize heterogeneous and sequential patterns of renewable energy droughts. Single-threshold analyses, which are prevalent in the literature, are not capable of detecting such patterns and lead to an incomplete characterization of extreme droughts.
@@
Based on a large data set and using a wide range of thresholds, we demonstrate that the complementarity of wind and solar power in Europe effectively reduces both the frequency and maximum duration of VRE droughts. Analyzing isolated countries, we find that combining solar and wind power leverages a portfolio effect, decreasing the maximum drought duration compared to single renewable technologies by on average 64% (PV), 52% (onshore wind), or 47% (offshore wind) in our European case study. This effect is particularly significant for PV, as wind power compensates for diurnal and seasonal shortages of solar energy. We further show that an unconstrained geographical balancing between European countries gives rise to a substantial balancing effect, which in our case study shortens the longest PV, onshore wind, offshore wind, or technology portfolio droughts by 1%, 46%, or 34%, or 65% on average. This effect is driven by an imperfect spatio-temporal correlation of extreme drought periods, especially concerning wind droughts.
@@
To overcome the challenges of single-threshold analyses, we introduce a drought mass metric that enables an integrated assessment of the duration and severity of events. It reveals that extreme drought events of varying severity may emerge sequentially within contiguous low-availability periods lasting several weeks or months. These events affect multiple European countries simultaneously, yet with varying intensity. In most countries and in a perfectly interconnected Europe, the largest drought events occur in winter. In wind-dominated systems, however, the most severe events may occur during summer. We show that compound winter events, characterized by extreme technology portfolio droughts in combination with peak electricity demand, determine the major discharging period of long-duration electricity storage in a fully renewable European energy system.
@@
The severity of such events varies substantially across years and countries, enabling geographical balancing to mitigate the need for system flexibility. Using the drought mass metric and drawing on renewable capacity assumptions from a policy-relevant scenario, we find that the most severe renewable technology portfolio drought observable in the data occurred in the winter of 1996/97, assuming a perfectly interconnected, fully renewable European energy system. This European super drought event lasted 55 days. The maximum droughts in individual, isolated countries are much longer, e.g., 109 days in Germany. This shows that even during the most pronounced European drought event, geographical balancing can be leveraged.
@@
Our multi-threshold analysis facilitates a comprehensive characterization of renewable drought events in Europe, solely based on VRE availability time series. Our approach, and particularly the drought mass metric, is not prone to artifacts that may arise in drought analyses relying on energy system modeling such as the duration of the planning horizon or technology cost assumptions. These may substantially affect optimal long-duration storage investments [dowling_role_2020]. Yet, methodological challenges remain. Drought analyses purely based on renewable availability time series cannot take into account real-world interconnection levels, which are between our extreme assumptions in the island and copperplate scenarios. Drought characteristics of scenarios with cross-regional exchange are thus likely between these corner solutions. For instance, the duration of the European super drought is likely to last longer than the 55 days identified for the copperplate scenario, but shorter than identified here for isolated countries (Figure 6), depending on future interconnection capacities between countries.
@@
Due to computational limits, energy system models are often solved for a limited number of weather years, or even for only a single weather year. While substantial inter-annual variation in optimal energy model outcomes has been documented in the literature [pfenninger_long-term_2016, staffell_using_2016, pfenninger_dealing_2017, zeyringer_designing_2018, collins_impacts_2018, schlott_impact_2018, kaspar_climatological_2019, hill_effects_2021, tong_geophysical_2021, grochowicz_intersecting_2023, gunn_spatial_2023, ruggles_planning_2024], our analysis shows that extreme portfolio droughts also vary significantly across different years and regions. This likely aggravates inter-annual variations of energy model outcomes, particularly in terms of long-duration system flexibility needs. Accordingly, it appears desirable to corroborate the findings and conclusions of previous model-based energy system studies that are based on a small number of weather years or even only a single one and do not include extreme VRE droughts. For example, the European Ten Year Network Development Plan 2022 draws on the single weather year 2009 [entso-e_tyndp_2022]. Long-term energy scenarios developed for the European Commission generally lack transparency about actual weather years used, but appear to be based on a very limited number of weather years [european_commission_directorate_general_for_energy_eu_2021]. The same holds true for influential climate neutrality scenarios for Germany [bdi_klimapfade_2021, fraunhofer_isi_langfristszenarien_2024, prognos_klimaneutrales_2021, luderer_deutschland_2021]. Our analysis shows that the winter of 1996/97 including the pan-European super drought is particularly relevant for weather-resilient planning of the European energy system. It further indicates relevant weather years for single-country analysis.
@@
Renewable drought analysis based on renewable availability time series, as conducted in this paper, is not intended for precisely quantifying long-duration storage requirements. Instead, it complements energy system modeling, which is required for evaluating flexibility needs in renewable energy systems for the reasons discussed above. Insights on extreme droughts can support the selection of relevant weather years for computationally intensive multi-carrier and multi-sectoral energy system models. Additionally, such insights can help explaining the outcomes of energy models, particularly concerning long-duration flexibility needs.
We see several promising avenues for future research. We note that a wide range of different datasets of renewable availability are used for energy system modeling, with sometimes conflicting model outcomes and policy recommendations [kies_critical_2021]. Assessing the sensitivity of VRE drought characteristics to such datasets could provide valuable feedback to the meteorological research community that generates and continuously improves such datasets [craig_overcoming_2022]. Further, the temporal extent of most data sets is limited to a few decades. Statistically robust findings for low-probability VRE drought events with high impact are likely to require longer records of weather data, which also might involve a need for synthetic datasets, e.g., applied in Gangopadhyay et al. [gangopadhyay_role_2022]. Moreover, drought analyses for modeled future weather data that incorporate the effects of global climate change would be desirable. Finally, characterizing renewable energy droughts in other world regions and comparing these with our findings for Europe would be of interest. While parameters like the cut-off threshold of the drought mass metric ought to be reviewed, we argue that our multi-threshold framework is particularly useful for respective analyses in world regions where renewable availability is characterized by diurnal or seasonal variability, e.g., related to regional weather phenomena such as monsoon events.
@@ 4.1 Variable renewable energy drought definition and identification
Our VRE drought analysis is based on hourly VRE availability factor time series. An hourly availability factor, also referred to as capacity factor, ranges between 0 and 1 and indicates the generation potential (in MWh) of a VRE source in a specific region and hour normalized by its capacity (in MW). We use a new open-source method that applies a variable-duration mean below threshold (VMBT) algorithm for drought identification. This method searches for periods with a moving average below a given drought qualification threshold by iteratively decreasing the event duration. In each iteration, the algorithm sets the averaging interval to the respective event duration, starting with very long-lasting events and iteratively continuing to such that last only a few hours. A moving average below the drought threshold identifies a drought event. It is excluded from subsequent iterations, in which the averaging interval decreases further and additional (shorter) events are identified. The algorithm is introduced and discussed in Kittel & Schill [kittel_measuring_2024].
@@
This iterative procedure overcomes shortcomings of previous research [kittel_measuring_2024]. It pools adjacent periods that independently may not qualify as VRE drought, identifies unique events with the longest duration possible, avoids double counting as well as overlaps with adjacent events, and ensures that the full temporal extent of drought periods is captured [kittel_measuring_2024]. To cover the full spectrum of VRE droughts, we parameterize the VMBT algorithm to iteratively search for periods ranging in duration from two years to one hour in descending order. The code and input data of the drought analysis tool are freely available on GitLab at https://gitlab.com/diw-evu/variable_renewable_energy_droughts_analyzer.
While threshold-based renewable drought identification methods are commonly applied in different bodies of the academic literature [kittel_measuring_2024], previous threshold-based analyses of renewable droughts across multiple regions, technologies, or technology portfolios often apply uniform, exogenously defined thresholds e.g., [leahy_persistence_2013, cannon_using_2015, patlakas_low_2017, kaspar_climatological_2019, ohlendorf_frequency_2020, kies_critical_2021, potisomporn_extreme_2024, potisomporn_extreme_2024, abdelaziz_assessing_2024, mockert_meteorological_2023, mayer_probabilistic_2023, ohba_climatology_2022]). Such approaches implicitly assume that one unit of generation capacity has a uniform generation potential across all region-technology settings. However, substantial differences in annual generation potentials may prevail due to varying meteorological conditions, leading to comparability issues in the resulting drought characteristics. To address this, threshold-based drought analyses should account for cross-regional and cross-technological variations in renewable generation potential when defining drought thresholds. Thresholds scaled relative to the mean availability of each setting (termed here as “relative thresholds”), reflect these variations [kittel_measuring_2024].
Additionally, as highlighted in the literature review, no consensus has yet been established regarding the selection of concrete threshold values. On the contrary, they often appear arbitrarily chosen or lack clear justification [kittel_measuring_2024].
@@
To address both the challenge of comparability across regions and technologies and appropriate threshold value selection, we apply a wide range of relative drought thresholds, each scaled relative to the multi-year mean availability of the respective time series, i.e., each single technologies or different VRE portfolios across countries. Threshold values $\tau$ range from 10% to 100% of the mean availability factor over all investigated years, increasing in 5% increments. This broad range allows us to capture a variety of renewable drought events. Very low thresholds capture potentially brief but very severe drought events with near-zero renewable availability (e.g., $\tau=0.1$), whereas high thresholds (e.g., $\tau=1.0$) are more likely to identify very long-lasting events or even full weather years with below-average renewable availability. Figure SI.1 illustrates the relationship between these relative thresholds and absolute mean availability factors.
@@
We use country-level VRE availability time series from the Pan-European Climate Database provided by the European Network of Transmission System Operators for Electricity [de_felice_entso-e_2022], comprising 38 weather years from 1982 to 2019. These data have been derived from reanalysis data, which has been converted to renewable availability by the Transmission System Operators. The database has been used for policy-relevant strategic reports, such as the Ten Year Network Development Plan (TYNDP) 2022 or the European Resource Adequacy Assessment 2021. We use exogenous renewable capacity assumptions to generate the capacity-weighted VRE portfolio time series as explained in the next paragraph. These assumptions stem from the TYNDP 2022 (scenario Distributed Energy), which optimized these portfolios for a climate-neutral European power sector in 2050. For Germany, we update the capacity targets according to the latest government targets for 2045 [bundesamt_fur_justiz_gesetz_2024], in which Germany aims to achieve climate-neutrality.
@@
4.2 Drought mass: a multi-threshold metric to identify storage-defining drought events
Here, we introduce a drought mass metric to identify storage-defining drought events in a renewable energy systems. This metric identifies and ranks extreme events by integrating both drought duration and severity across a wide range of thresholds. As discussed in Section 2.3, very high relative thresholds close to $\tau=1.0$ are not meaningful, as they tend to capture very extended periods or even whole weather years with below-average renewable availability, instead of unusual drought events. For the drought mass metric computation, we therefore limit our analysis to drought events identified at thresholds $\tau\in[0.1,0.15,...,0.75]$, a set with cardinality $c=14$. For each threshold $\tau$, we first generate binary time series for each year in the data, which assign the value $1$ to hours that qualify as drought and 00 otherwise. These time series are then paired into consecutive two-years periods (e.g., 1982-1983, 1983-1984, 1984-1985, etc.), resulting in vectors with a magnitude of $T=17520$ hourly time steps per pair. This results in 37 year-pairs and accordingly allows analyzing 37 complete winter droughts. Note that the investigated data abstract from leap days. Next, we stack binary threshold-specific vectors to form a matrix of dimensions $T\times c$, where each row corresponds to a specific hour and each column to a threshold. Within each two-year period, we search for contiguous drought events identified at $\tau=0.75$, starting at hour $k$ and ending at hour $l$. For each of these intervals $[k,l]$, we compute the total number of drought hours across considered thresholds, i.e., we sum across all $c$ columns. The event with the highest cumulative multi-threshold score is then selected as the most extreme drought event for each weather year pair.
The drought mass metric equally weighs drought events identified across all considered thresholds. We chose a cut-off threshold $\tau=0.75$, which defines the event duration. Within an contiguous event, all thresholds $\tau\leq 0.75$ equally contribute to an event’s overall score. We tested a wide range of alternative cut-offs and weighting schemes to identify the most effective drought mass design. The best alignment between events with the highest winter drought mass scores and major storage discharge events was achieved at a cut-off threshold of $\tau=0.75$ and a simple approach of equal weighting across all considered thresholds.
@@ 4.3 Power sector modeling and scenarios
We use the open-source power sector model Dispatch and Investment Evaluation Tool with Endogenous Renewables (DIETER) to analyze the interaction between VRE droughts and long-duration storage needs in a fully renewable European power sector. DIETER is a linear optimization model that determines least-cost capacity and dispatch decisions based on an hourly resolution [zerrahn2017, gaete-morales_dieterpy_2021]. Different versions of the model have been used to study various aspects of VRE integration and their interaction with other flexibility options or sector coupling technologies [zerrahn2018, schill_long-run_2018, schill_electricity_2020, schill_flexible_2020, kittel_renewable_2022, roth_geographical_2023, kirchem_power_2023, gong_bidirectional_2023, gaete-morales_power_2024, roth_power_2024]. Here, we use a model version that includes 33 European countries (EU27, the United Kingdom, Norway, Switzerland, and the Western Balkans). As very long-lasting VRE droughts may span across the turn of a calendar year, we extend DIETER’s planning horizon to two full years. For transparency and reproducibility, we provide the model code, the input data, and a manual in a public repository under permissive licenses available on GitLab at https://gitlab.com/diw-evu/projects/quantifying_the_dunkelflaute_energy_system_analysis.
@@ Acknowledgments
We thank the entire research group “Transformation of the Energy Economy” at the German Institute for Economic Research (DIW Berlin) for valuable inputs and discussions, as well as conference participants of the EGU General Assembly 2022 and 2025, the International Energy Workshop 2022, the Conference on Climate, Weather and Carbon Risk in Energy and Finance 2022, the Next Generation Energy and Climate Workshop 2022, the International Conference Energy & Meteorology 2023, and the ENERDAY conference 2024 for valuable comments on earlier drafts. We acknowledge research grants by the Einstein Foundation (grant no. A-2020-612) and by the German Federal Ministry of Education and Research via the “Ariadne” projects (Fkz 03SFK5NO & 03SFK5NO-2).
@@
SI.2 Additional illustrations of identified drought patterns
Figure SI.2 illustrates identified droughts lasting longer than one week or more, filtering out briefer events. This highlights very long-lasting below-average wind periods that may encompass multiple drought events of varying severity and solar seasonality. Combined in a portfolio, these periods are less severe (portfolio effect) and are further mitigated when assuming unconstrained geographical balancing (balancing effect). Figures SI.3SI.4, and SI.5 show identified drought patterns lasting at least one day for the years 2011/12, 1982/83, and 2013/14.
Figure SI.2: Identified drought patterns in 1996 and 1997 across all employed thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one country. To illustrate very persistent shortage situations, only droughts lasting longer than one week are displayed.
Figure SI.3: Identified drought patterns in 2011 and 2012 across all employed thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one country. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. In Spain, low PV seasonality is accompanied by wind droughts, translating into portfolio droughts arising across the turn of the year 2011/12. In contrast, due to the exceptionally high wind availability in North Europe, no portfolio droughts occur across the turn of the year in these countries.
Figure SI.4: Identified drought patterns in 1982 and 1983 across all employed thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one country. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. Except for Spain, no portfolio droughts occur at the beginning of a calendar year due to the absence of wind droughts in Northern Europe.
Figure SI.5: Identified drought patterns in 2013 and 2014 across all employed thresholds. For each technology-specific panel, a horizontal band indicates drought occurrences for the color-coded threshold of one country. To illustrate persistent patterns, only droughts lasting longer than one day are displayed. Brief wind droughts are infrequent in winter but longer-lasting, more severe, and more frequent in summer. Therefore, less severe portfolio droughts occur in the winter of 2013/14.
SI.3 Additional illustrations of frequency-duration distributions
@@
Figure SI.6 illustrates the cumulative frequency-duration distributions for droughts lasting up to a full year for Germany, Spain, and the European copperplate scenario. Only higher thresholds identify droughts lasting longer than a few weeks. For thresholds $\tau<1$, portfolio droughts are generally briefer than single-technology droughts. To raise complementary insights on seasonality, Figure SI.7 shows seasonally differentiated distributions for the copperplate scenario. In general, longer droughts are more frequent in winter than in summer. Brief PV droughts are more frequent in summer and are detected by higher thresholds.
@@
Figure SI.6: Example of cumulative frequency-duration distributions of drought events across all investigated thresholds $\tau\in[0.1,1]$ that may last up to one full year, sorting the frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific frequency.
@@
Figure SI.7: Example of cumulative frequency-duration distributions of drought events across all investigated thresholds $\tau\in[0.1,1]$ under the assumption of unconstrained geographical balancing, sorting the frequencies of all events that are at least as long as a given duration. White space indicates the absence of droughts for given thresholds in the data. The contour lines represent the threshold-specific frequency. For illustration, the distributions are truncated at 360 hours, i.e., they show events with a maximum duration of just above two weeks. Frequencies of events lasting at least two days, one week, and a fortnight are marked for a relative threshold $\tau=0.75$.
SI.4 Return periods: most extreme events occur rarely
@@
Return periods of VRE droughts are given by the reciprocal of yearly frequencies. The return period-duration distributions represent the right-hand side of the event distribution and indicate the longest event duration that can be expected to reoccur after a given number of years (Figure SI.8). Note that the return period monotonically increases and the maximum return period is limited by the 38 years of data that we investigate.
@@
For each threshold, the maximum event duration increases with higher return periods (compare increasing contour lines and marked data points in Figure SI.8). Additionally, the return period-duration distributions are highly sensitive to the underlying threshold. For any given return period, higher thresholds lead to events with longer maximum durations, plateauing at 365 days for very high thresholds. Note that such high durations indicate below-average wind and solar years and should not be mistaken for actual drought events. The return period-duration distributions vary significantly between technologies and countries. For example, the maximum offshore wind drought for a relative threshold of $\tau=0.75$ in Germany that returns every 20 years lasts 198 days and is much shorter than the corresponding PV (289 days) and onshore wind droughts (297 days). By comparison, in Spain, such 20-year return period droughts are shorter for PV (216 days) and onshore wind (245 days) but longer for offshore wind (211 days).
@@
When combined in a technology portfolio, the maximum duration decreases substantially for each return period in both Germany and Spain and also in the European copperplate, which is another instance of the portfolio effect described in the main body of this paper. The balancing effect is also visible: with perfect interconnection, the maximum drought duration further decreases due to a limited temporal correlation of wind droughts across Europe. In the case of solar PV, higher solar availability in South Europe during winter balances lower availability in Northern Europe.
@@
Figure SI.8: Example of return period-duration distributions of rare drought events with an average annual frequency below 1 across all investigated thresholds $\tau\in[0.1,1]$. The contour lines along the return period axis represent the threshold-specific return period. The maximum duration of an event returning on average every 10, 20, and 30 years are marked for a threshold of $\tau=0.75$.
SI.5 Additional illustrations of maximum drought durations
@@
Figure SI.9 shows the longest droughts obtained from the data for each year and threshold for Germany, Spain, and the European copperplate scenario. Figures SI.10 and SI.11 show the maximum duration of single drought events across all years for each country and threshold. Figure SI.12 illustrates the inter-annual variability of the maximum drought duration for each threshold and technology (portfolio).
Figure SI.9: Examples of most extreme duration of single drought events for each year and across all investigated thresholds $\tau\in[0.1,1]$. The contour lines represent threshold-specific maximum duration. White space indicates the absence of droughts for given thresholds in the data. The events with the highest and lowest duration across all years are marked for a threshold $\tau=0.75$. Arrows indicate values that are hidden in valleys of the distribution plane.
@@
Figure SI.10: Most extreme duration of single drought event of all years for each country and across all investigated thresholds $\tau\in[0.1,1]$. The latter are sorted in descending order from left to right according to the longest duration for a threshold $\tau=0.75$.